A cylindrical tube, open at both ends, has fundamental frequency f in air. The tube is dipped vertically in water so that half of it is in water. The fundamental frequency of the air column is now
(a) f/2
(b) 3f/2
(c) f
(d) 3f

Grade:Upto college level

2 Answers

Arun
3 years ago
f=2lv

Now, it will act as one end open and one end closed.

So, f0=2lv=42lv=2lv=
 
Hope it helps
 
Thanjs and regqrds
Vikas TU
3 years ago
Case I Here λ/2=l
∴λ=2l
Now, v=f×λ
∴f=v/λ=v/2l ….(i)
Case II Here λ'/4=l/2
∴λ'=2l
Now, v=f'×λ'
∴f'=v/λ'=v/2l=f.